OpenStudy (anonymous):

Use the properties of log to write each expression as a single log with coficient of 1: ln(3) -2ln(4) +ln(8) got this wrong on my quiz and need to figure out how to do for mid-term Please show work..I wrote .41

OpenStudy (anonymous):

Here are some of the rules for ln's (natural logarithms)(or logs in general). log(xy) = log x + log y log(x/y) = logx - log y log x^p = p log x See if you can use them to answer the question...

OpenStudy (anonymous):

I'll start u off, using the third rule, 2ln4 = ln 4^2

OpenStudy (anonymous):

Ln 3 - 2ln4 + ln 8 => Ln 3 - Ln (4^2) + Ln 8 => [Ln3+Ln8] - Ln16 =>Ln(3*8)- Ln 16 => Ln 24 - Ln 16 => Ln (24/16) => Ln(3/2).. < *Answer*

OpenStudy (anonymous):

I deserved a medal there, though! For that elaborate answer! :P